Mary optimizes her satisfaction when she consumes 10 books (B) and 5 apples (A) each month. Mary's monthly income is $ 50 and she currently spends $ 10 on apples. Write Mary's budget line in slope intercept form with books measured on the vertical axis.


1
Expert's answer
2021-01-13T10:24:03-0500

Pa = $10/5 = $2.

"Pb = (50 - 10)\/10 = 4" $.

Mary's budget line is 2A + 4B = 50 or B = 12.5 - 0.5A.